Solve the system of equations and enter the solution as an ordered pair.
2x-5y=-6
8x+3y=68

Answer by texttutoring(324)

2x-5y=-6
8x+3y=68

Multiply the first equation by 4 (so that both equations will have 8x terms) and then subtract:
8x-20y = -24
8x +3y = 68
------------
0 -23y = -92

Now divide both sides by -23:

y = 4

Find x by plugging y=4 into either equation:

8x +3(4) = 68
8x = 68 - 12
8x = 56
x = 7

So the answer is ordered pair is (7,4)